### v1

# Personality Pairing Improves Human-AI Collaboration

Abstract
Here we ask how AI agent “personalities” interact with human personalities, and other traits, to shape human-AI collaboration, productivity and performance. To estimate these relationships, we conducted a large-scale preregistered randomized experiment that paired 1,258 participants with AI agents that were prompted to exhibit varying levels of the Big Five personality traits. These human-AI teams produced 7,266 display ads for a real think tank, and the quality of these ads was evaluated by 1,995 independent human raters as well as in a field experiment conducted on X, which generated nearly 5 million impressions. We found, first, that personality pairing impacted teamwork quality. For example, neurotic AI improved teamwork for agreeable humans but impaired it for conscientious humans. Second, we found productivity effects of personality pairing and a “productivity-performance trade-off" in which certain pairings (e.g., agreeable human with neurotic AI) produced fewer ads but of higher quality. Third, personality pairing influenced ad quality and performance. For example, quality improved when open humans were paired with conscientious AI and when conscientious humans were paired with disagreeable AI. Some of these pairing effects were “jagged” in that they varied across text and visual tasks. For example open humans produced higher quality images but lower quality text when paired with agreeable AI. Pairing effects were also present in other human traits, like country of origin. For example, extroverted AI improved quality for Latin American workers, but degraded quality for East Asian workers. These findings demonstrate that human-AI personality alignment significantly improves collaboration, productivity, and performance and lay a foundation for future research on improving human-AI collaboration through AI personalization.
Introduction
Artificial intelligence (AI) tools have gained significant attention for their ability to increase productivity and performance across a wide range of tasks. For example, chatbots powered by large language models (LLMs) reduced the time required for mid-level writing tasks by 40% while improving quality by 18% (noy2023experimental). AI-enhanced resumes increased job hiring rates by 8% (wiles2023writing). In customer support settings, AI assistance increased productivity by an average of 14% (brynjolfsson2023genai). But these benefits also exhibited considerable heterogeneity, varying across tasks (dellacqua2023navigating, ju2025collaborating), worker seniority (wang2023friend), gender (dellacqua2023navigating, humlum2024adoption), and skill levels (noy2023experimental, brynjolfsson2023genai, otis2024uneven). In clinical settings, AI assistance reduced reading time while maintaining diagnostic accuracy, but effects varied significantly across physicians and clinical contexts (chen2024impact, yu2024heterogeneity, agarwal2023combining). In financial services and customer support, human involvement in AI-based services boosted consumer confidence and material outcomes, though in other contexts AI reduced quality due to sycophancy and over-reliance (yang2025my, vaccaro2024combinations, sergeyuk2025human). Unfortunately, experiments documenting these productivity and performance gains have generally provided participants with identical AI tools, leaving open the possibility that at least some of the observed variation in performance stems from a mismatch between AI and individual human characteristics. Performance gains may therefore increase and simultaneously vary less by age, gender, or skill level, if AI is personalized for each human collaborator.
Such personalization may now be possible as foundation model companies have recently begun developing, refining and deploying AI personalities to improve human-AI interaction. For example, OpenAI introduced personality customization in models like GPT-5, offering Default, Cynic, Robot, Listener, and Nerd personalities (openai2025customizing, openai2025sycophancy). Similarly, Anthropic has begun to measure undesirable personality shifts during training, such as sycophancy, a tendency for AI to overly align with user preferences (anthropic2024character, sharma2025sycophancy), or other “evil” traits, and introduced novel training methods to mitigate such personality shifts in their models. These advances enable AI agents to be tailored to users, potentially improving human-AI collaboration and ushering in a new era of AI personalization.
Despite the development of AI personalities, however, no prior research has rigorously and experimentally examined their causal effects on human-AI collaboration. Such experiments are now critical to understanding whether and how tailored AI personalities may impact teamwork, productivity, and performance as the lack experimental evidence on AI personalization hinders the optimization of AI systems for collaborative tasks and undermines our understanding of the detrimental effects of mismatched AI personalities on, for example, AI psychoses.
In human teams, team composition is a critical factor that promotes trust, coordination, and teamwork (zhang2023teammate, hemmer2023, woolley2023collective). Humans exhibit variation in demographic, psychological, and behavioral traits known to impact team performance (kichuk1997, peeters2006, woolley2010ci). Theories of person-team fit suggest that when team members’ personalities are compatible, collaboration becomes more effective and leads to better results (barrick1998gma, bell2007meta). As AI has evolved from a passive tool to an active collaborator (makarius2020rising, fugener2022cognitive, anthony2023, collins2024building, ju2025collaborating), tools like Cursor or Claude Code now act on behalf of and co-create alongside human users, as team members, in real time. Like humans, these AI agents also possess characteristics that influence teamwork, technology acceptance, and even psychosocial outcomes (li2025acceptance, tully2025lower, fang2025psychosocial).
Unfortunately, although human and AI traits shape human-AI teamwork, their interaction remains understudied (krakowski2025human). With AI agents increasingly operating as social actors in collaborative workflows (anthony2023, gelbrich2025), it is crucial to understand whether the principles of personality alignment apply to human-AI teams and whether AI agents can be designed to exhibit personalities that complement those of their human counterparts. To address these gaps in our understanding of human-AI collaboration, we examined how AI personalities interact with human characteristics to shape teamwork, productivity, and performance.
Our study leverages the unique capabilities of AI agents to vary their individual personalities (jiang2023evaluating, chen2025persona), allowing us to systematically investigate how human-AI personality alignment influences the productivity and performance of human-AI teams. Systematic prompt variation can effectively shape AI agent behavior (schulhoff2024prompt), and seminal work on OpenAI’s ChatGPT architecture demonstrates that prompting enables control of agent behavior without task-specific fine-tuning (brown2020language). Advanced prompting methods, such as chain-of-thought prompting, have further demonstrated significant improvements in task performance by guiding a model’s reasoning processes (wei2023chainofthought) and “persona vectors" have been shown to successfully monitor and control personality shifts in AI agents (chen2025persona). These methods enable us to intentionally design AI agents with particular social and behavioral characteristics, which can be tailored to complement human collaborators. Randomized controlled trials (RCTs) are now needed to evaluate the effects of such AI personalities on human-AI collaboration and task outcomes to understand how AI agents can be personalized for effective teamwork (jakesh2023opinionated).
Investigating human-AI collaboration with agentic tools is challenging in practice. These tools must be seamlessly embedded in workspaces, have full contextual awareness, and be capable of taking actions on behalf of users. Developing such sophisticated tools is a complex and costly endeavor, and rigorous research in this area requires multiple treatment arms to test interactions with individual covariates, necessitating large sample sizes to ensure statistical power. Downstream performance measures are also crucial to evaluating the effectiveness of real collaborations in the field, not just in the lab. Our study addresses all of these requirements, providing a comprehensive framework for investigating the real-world economic impact of AI personalization.
Experiment Design
To investigate how personality alignment between humans and AI agents influences collaboration, we conducted a large-scale experiment involving 1,258 participants representative of the U.S. population using Pairit, an experimental platform we developed for extensible human-AI collaboration (ju2025collaborating). After taking a 10-item pre-task survey to measure their Big Five personality traits, participants, who were recruited via Prolific from October 15 to 18, 2024, were randomly paired with multi-modal AI agents, which had full context of the workspace and were induced to independently exhibit high or low levels of the Big Five personality traits using P2 prompting (jiang2023evaluating). These human-AI teams then worked together to create display ads marketing the year-end report of a real think tank within a 40-minute session, using real-time chat and synchronized text and image editing tools. Teams were provided with image assets by the think tank and could separately call Dall-E 3 to generate new images for the campaigns. As described in the Materials and Methods, the AI agents were fully multi-modal, had access to the Internet, could manipulate the images, the image carousel and the text of the ads, text with their human collaborators and could also call Dall-E 3 to generate new images. This setup allowed us to assess the causal effects of personality alignment on collaboration outcomes. The randomization of AI personalities, such as high or low conscientiousness, openness or extroversion, enabled us isolate the causal effects of human-AI personality pairing on the collaboration, productivity and performance of human-AI teams.
Figure 1: The experimental design for investigating human-AI collaboration. (A) Participants are randomized into collaborating with either a human partner or an AI agent. (B) AI agents are randomly assigned prompts to induce low or high levels of the Big Five personality traits. (C) Pairs collaborate in a real-time collaborative workspace to create display ads.
Importantly, we approach personality prompts as neither about memory, such as recalling past chat interactions, nor about aligning LLM responses to human preferences which often leads to AI sycophancy (chhikara2025mem0, sharma2025sycophancy). Instead, we randomize AI personalities to learn the causal effects of various personality pairings on human-AI team performance. This approach parallels findings in human-robot interaction, where robots’ social communication and appearance have been shown to influence team performance (jung2013backchanneling, jung2018theories).
To assess the quality of the display ads produced by the human-AI teams, we conducted a separate evaluation study using a second independent sample of 1,198 new participants recruited from Prolific. These raters were based in the US and selected to ensure representation across gender, age, and ethnicity. Each rater evaluated a random sample of 40 ads, with the sampling design guaranteeing that every ad received at least three independent ratings. Raters assessed each ad on three dimensions using 7-point Likert scales: text quality (“The text is present, clear, relevant, and engaging”), image quality (“The image is visually appealing”), and likelihood of clicking the ad (“I am likely to click on this ad”). The ads were presented as realistic mockups, incorporating standard display ad elements such as images, copy, call-to-action buttons, and social media-style interaction options (e.g., Like, Comment, Share; see Supporting Information [SI] Section C).
To evaluate the real-world performance of the display ads, we then conducted a pre-registered field experiment on the X platform. We sampled 2,000 ads from the total 7,266 ads using stratified sampling based on predicted click likelihoods implied by the human ratings to ensure sampling across the ad quality distribution. These ads were then run in 400 geographically non-overlapping campaigns, each with a five-ad split test (braun2024abtesting, braun2024adplatforms). Key performance metrics included the click-through rate (CTR) and the cost-per-click (CPC) measured on X, as well as the view-through rate (VTR; fraction of document viewed) and the view-through duration (VTD; seconds) tracked via unique DocSend links for each ad (details of the field experiment are provided in the Materials and Methods).
Results
Our findings revealed that personality pairings significantly affected collaborative outcomes including teamwork quality, productivity, ad quality and ad performance. For example, humans high in conscientiousness created ad text of greater quality when paired with conscientious AI agents, while extroverted humans paired with conscientious AI produced lower quality text, lower quality images, and ads with lower click likelihood. Some pairings demonstrated uneven “jagged” performance patterns, with effects varying sharply between the textual and visual components of the task. We further observed effects on productivity, with agreeable humans, for example, producing more ads when paired with open AI. Additionally, certain combinations yielded a performance-productivity trade-off, such as agreeable humans with neurotic AI generating fewer ads, but ads of higher quality. Overall, these results underscore the substantial role of personality alignment in human-AI teamwork and lay a foundation for future research on AI personalization to improve human-AI productivity and collaborative performance. In particular, we found specific effects on teamwork, productivity and output quality.
Teamwork
First, we found personality pairings significantly influenced teamwork quality. To evaluate these effects, participants completed a 35-item teamwork quality survey after the task, which assessed six facets of collaboration: communication, coordination, balance of member contributions, mutual support, effort, and cohesion (hoegl2001). We conducted a regression analysis on data produced by our experiments to identify the average treatment effects of AI personalities as well as the interaction effects of human and AI personality traits in combination and found statistically significant effects for human openness, conscientiousness, and agreeableness, as well as for AI neuroticism (Figure 2, see SI Section H).
Figure 2: Regression coefficients for human-AI personality interaction terms for teamwork quality (p<<0.05) (hoegl2001). Error bars show 95% confidence intervals. See SI Section H for the full regression tables.
In examining interaction effects, we found that for open humans, conscientious AI lowered the balance of member contributions, mutual support, effort, and cohesion. For conscientious humans, open AI improved balance and cohesion, and conscientious AI improved effort and cohesion. However, for conscientious humans, neurotic AI lowered communication, coordination, balance, and mutual support. Finally, for agreeable humans, neurotic AI improved communication, coordination, mutual support, and effort. However, these positive results reversed when agreeable humans were paired with open AI (for cohesion) and with conscientious AI (for coordination, balance, and cohesion). These effects revealed that personality pairings shape perceived teamwork dynamics in human-AI collaboration.
Ad quality
In human evaluations of ad quality, we found that personality pairing between humans and AI agents significantly influenced ad text quality (Figure 3) beyond the effects of AI personality alone (see SI Section H). The average treatment effects demonstrated that conscientious AI reduced, and extroverted AI improved, text quality on average. The interaction effects then demonstrated statistically and economically significant effects of human-AI personality pairing on ad quality. For both open and conscientious humans, conscientious AI increased text quality, but agreeable AI decreased text quality. For extroverted humans, extroverted AI increased text quality, but conscientious AI decreased text quality. For agreeable humans, neurotic AI increased text quality, but extroverted AI decreased text quality. These results indicate that aligning AI personality traits with human traits impacts the effectiveness of collaborative ad text production.
Figure 3: Regression coefficients for human-AI personality interaction terms on ad text quality, image quality, and click likelihood (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.
For image quality, we observed a different set of significant interaction effects based on human-AI personality pairing. For open humans, agreeable AI increased image quality while, for conscientious humans, open AI increased image quality. For extroverts and agreeable participants, being paired with conscientious and open AI, respectively, decreased image quality.
For click likelihood ratings, which reflect the perceived engagement potential of the ads, we observed no significant effects for humans with high openness or agreeableness. For conscientious humans, agreeable AI decreased click likelihood, while for extroverted and neurotic humans, conscientious AI decreased click likelihood.
Across the three measures of ad quality (text quality, image quality and perceived click likelihood), personality pairings between humans and AI agents consistently influenced outcomes. For extroverts, conscientious AI decreased quality across all measures. For conscientious humans, agreeable AI decreased text quality and click likelihood. Some pairings had significant and opposite effects, indicating a “jagged edge” of human-AI competence across tasks and personality pairings (dellacqua2023navigating). For example, for open humans, agreeable AI improved image quality but worsened text quality. Collectively, these patterns suggest that strategic personality pairings can optimize human-AI collaboration outcomes, with some generalizable and some task-specific personality pairings.
In addition to participant personality traits, we found that participant demographics, including previous AI experience, employment status, and country of origin, moderated the effects of AI personality traits on ad quality. For individuals with previous AI experience, extroverted AI hurt overall performance, while neurotic AI improved performance (see SI Sections G).
Although all participants were based in and representative of the demographic composition of the U.S., their diverse countries of birth allowed us to explore possible cultural and regional differences in collaborative outcomes (Figure 4, see SI Section H). Participants from Western regions, the Middle East & South Asia produced significantly lower quality images when paired with agreeable AI, while participants from East Asia produced images of significantly lower quality when paired with extroverted and agreeable AI. Participants from Latin America produced significantly higher quality ad text when paired with extroverted AI and significantly lower quality ad text when paired with conscientious, agreeable, and neurotic AI. For participants from all regions except East Asia, agreeable AI significantly lowered click likelihood. For those from Latin America, extroverted AI improved click likelihood, while conscientious AI lowered click likelihood. For those from the Middle East & South Asia, agreeable AI lowered click likelihood.
Interestingly, we observed that agreeable AI consistently undermined performance across all measures, reducing image quality (for Western, East Asian, Middle Eastern & South Asian workers), text quality (for Latin American workers), and click likelihood (for Western, Latin American, Middle Eastern & South Asian workers). Taken together, these patterns suggest individual characteristics like culture mediate how AI personality traits, especially agreeable or sycophantic AI (sharma2025sycophancy), affect collaboration and suggest the potential for future research investigating the strategic pairing of AI personalities with various human traits.
Figure 4: Regression coefficients for interaction terms of a participants’ country of birth and AI personality on ad quality measures (p<<0.05). Error bars show 95% confidence intervals. See SI Section H for the full regression table.
Ad performance
We also evaluated the real world performance of the ads produced by our human-AI teams in a large-scale field experiment on X. Our field study evaluated traditional marketing metrics, including click-through rates (CTR) and cost-per-click (CPC), as well as downstream ad behaviors, including view-through rates (VTR) and view-through duration (VTD), using a unique DocSend link for each ad (Figure 5). VTR and VTD were measured to assess whether participants created click-bait ads irrelevant to the research report, which could elevate CTR but reduce VTR and VTD. In line with previous findings (ju2025collaborating), our mixed effects regression analysis revealed that higher quality ad text reduced CPC while higher quality images improved CTRs.
Figure 5: Regression coefficients for human-AI personality interaction terms and ad quality ratings on click-through rates (CTR; basis points) and cost-per-click (CPC; \$) and view-through rates (VTR; %) (p<<0.05). Higher values indicate better performance for CTR and VTR, whereas lower values indicate better performance for CPC. Error bars show 95% confidence intervals. See SI Section K for the full regression tables.
Personality pairings also significantly influenced ad performance. For conscientious humans, neurotic AI lowered CTRs, while for neurotic humans, neurotic AI increased CTR performance. Conscientious humans with neurotic AI and extroverted humans with agreeable AI both experienced significantly higher CPC. Open humans experienced lower VTRs when paired with conscientious AI, while conscientious humans produced ads with higher VTRs when paired with open AI. We found no statistically significant effects for VTD. Together, these results indicate that human-AI personality pairings influence ad performance both by shaping quality metrics and through effects on ad performance not captured by quality measures.
Productivity
In addition to evaluating the quality of the ads produced, we examined the effects of personality pairings on the productivity of human-AI teams. We defined productivity as the number of ad submissions completed by participants within a 40-minute task period. Since participants were given explicit monetary incentives to produce a greater number of ads and to produce higher quality ads, this definition offers a quantitative measure of output, and thus productivity, in the collaborative process of creating display ads.
Our analysis revealed several significant effects of personality pairing on productivity (Figure 6, see SI Section H). For extroverted humans, agreeable AI agents increased productivity, and for agreeable humans, open AI increased productivity while neurotic AI decreased productivity. These findings indicate that specific personality pairings influence productivity in human-AI collaboration. In particular, neuroticism mediates how agreeable, or sycophantic, AI impacts productivity. There were no significant average treatment effects for AI personality alone (see SI Section J).
Figure 6: Regression coefficients for human-AI personality interaction terms for productivity (i.e., the number of submissions; p<<0.05) for three human personality covariates, extroversion, agreeableness, and neuroticism. Error bars show 95% confidence intervals. See SI Section H for the full regression table.
Interestingly, some personality pairings introduced a trade-off between quality and productivity. Pairing agreeable humans with open AI increased productivity but decreased image quality. Pairing agreeable humans with neurotic AI had the opposite effect—they produced fewer ads but with higher text quality. These results suggest a trade-off in human-AI collaboration whereby some personality pairings improve the productivity of human-AI teams in terms of the volume of output produced per unit time, while other pairings improve ad quality while degrading productivity.
Discussion
We provide the first experimental evidence that personality alignment improves human-AI collaboration and team performance. Our work highlights the need to treat AI agents not just as tools, but as teammates whose traits can be intentionally engineered for collaborative success. Recently, others have tailored human approaches to working with AI, for example by shaping how salespeople interact with AI under different decision-making authority, training, and incentives conditions, to improve sales performance (krakowski2025human). To our knowledge, ours is the first study to measure the productivity, performance and teamwork effects of AI personalization in a context in which AI personalities are directly and randomly manipulated. As AI agents become more prevalent in collaborative settings, designing them with personality traits that complement their human counterparts could enable optimization of productivity and performance in real agentic workflows and unlock more effective and satisfying collaborations.
Our results extend the person-team fit literature to human-AI teams and reveal key patterns in human-AI collaboration dynamics (barrick1998gma, bell2007meta). For example, conscientious human-AI pairs demonstrated higher effort and cohesion, suggesting that shared conscientiousness creates a structured collaborative environment that promotes more effective interaction. In addition, agreeable AI, which is in essence sycophantic AI (sharma2025sycophancy), increased productivity for extroverted and neurotic people, but reduced all quality ratings when paired with conscientious people. This suggests that sycophantic or agreeable AI enables certain humans to produce more output per unit time, but erodes the quality of that output, perhaps because it pushes back less against human work and therefore fails to improve the work while letting a greater amount of lower quality work through to production. While agreeable AI can lead to over-reliance on AI in psychosocial contexts (fang2025psychosocial), our results suggest they can also be tuned to shape performance and productivity when tailored for individual human characteristics. This suggests that sycophantic AI could be managed toward productive outcomes when applied in the right contexts for the right people and avoided in the wrong contexts and for the wrong people. These findings highlight how different personality alignments activate distinct collaborative mechanisms, with some improving teamwork and others increasing productivity. The variation in productive outcomes associated with personality matching further demonstrates that effective human-AI collaboration requires strategic consideration of trait compatibility rather than a one-size-fits-all approach.
To make the practical implications of these results concrete, consider several vignettes that suggest particular pairing choices for conscientious, agreeable, extroverted and neurotic human workers. On one hand, conscientious people benefit from conscientious and open AI. On the other hand, agreeable people are more productive but produce lower quality work when working with open AI and produce higher quality work at a slower rate when working with neurotic AI. Extroverts benefit from extroverted AI, but produce lower quality work when collaborating with conscientious AI. Neurotic people perform better when paired with extroverted AI, but worse when working with conscientious AI. These vignettes highlight concrete design choices for matching AI personalities to workers to improve human-AI collaboration.
Although our study is the first randomized control trial of AI personalization in a real-world work task and provides the first empirical evidence that human-AI personality alignment can improve collaboration, productivity and performance in human-AI teams, it is not without limitations. First, the exact personality pairing effects discovered in our marketing context may not extend to other collaborative tasks, such as coding or analytics. Indeed, our analysis of the productivity and performance effects of human-AI personality alignment in a multi-modal ads creation task, which involved both visual ad images and textual ad copy, itself revealed that personality pairing effects can vary across tasks. Future research is therefore needed to measure pairing effects and their generalizability or jaggedness across tasks.
Second, our personality prompts were independently randomized and analyzed as interaction effects with human personalities. As such, we did not examine the performance effects of any optimization procedure or intentional pairing of human to AI personality traits. In this way, our results provide a rigorous, large-scale existence proof for the potential of optimized AI personalization, rather than exploring hypothesis driven, intentional or strategic personalization per se. While our extensive randomization allows us to analyze a wide range of diverse effects to efficiently generate hypotheses, this approach requires a large sample to have sufficient statistical power and becomes challenging to analyze in a theoretically driven manner. Future studies could focus on specific personality pairings to directly and efficiently answer theoretical hypotheses or evaluate optimized personalization, especially given the hypotheses generated from this study. Furthermore, while we employ a version of P2 prompting (jiang2023evaluating), new and developing methods may better steer AI toward desired personalities (chen2025persona), and these methods should also be investigated as avenues toward strategic AI personalization.
Third, our focus on the Big Five personality traits, though robust and widely used in human personality research, may overlook other personality traits potentially more relevant to human-AI dynamics. Our analyses indicate that demographics, like country of origin, moderate performance effects, suggesting broader heterogeneity in trait interactions that warrant further exploration. Future work could also investigate linguistic patterns to better understand traits that influence teamwork (wan2025personalized).
Fourth, the study’s single 40-minute session also limits understanding of how these relationships develop over time, such as over-reliance on AI (fang2025psychosocial). Overall, these considerations highlight avenues for future research, including in-depth textual and interaction analyses, to expand on this study’s contributions to human-AI collaboration.
Finally, our study says little about the psychosocial implications of AI personalization. Although our teamwork variables address self-assessed teamwork dynamics, we examine no measures of well-being or outcomes related to, for example, AI psychosis. As the field of AI personalization develops, it will be important to understand the psychological, social, and psychosocial implications of tailoring AI personalities to their human users and we encourage the development of that line work in future research.
Taken together, the effects shown here underscore the critical role of personality alignment in improving human-AI collaboration. By demonstrating that specific pairings can improve performance while others hinder it, we highlight the nuanced dynamics at play in human-AI teams. The observed trade-offs between quality and productivity further illustrate the complexity of these interactions. These insights pave the way for designing AI agents that are not only technically proficient but also socially and behaviorally adaptive, ensuring they complement their human counterparts effectively. Ultimately, our work demonstrates that personality-aligned AI has the potential to transform collaborative workflows and performance in human-AI teams.
Methods
This research was preregistered and approved by the MIT Committee on the Use of Humans as Experimental Subjects (see osf.io/jfzha and osf.io/95dhu).
The Pairit platform
The study was conducted on the Pairit platform (Figure 7). On the left panel of the platform, users can read the task, select images from a carousel, generate images using Dall-E, and write copy for the display ads. In the human-AI teams, the AI has full context of all elements on the platform, including screenshots of the ad images as they evolve over time during the collaboration. In all conditions, only the participants can submit ads. Additional details on the platform are in SI Section A-B.
Figure 7: The Pairit platform. The task workspace is on the left, and the messaging interface on the right. The participant chats with an AI agent which has full context of the user interface and can modify all elements on the task panel.
To enable the AI agents to act as effective collaborators on the Pairit platform, we provided them with full context of the task and collaboration by prompting each LLM call with all information on the user interface. The prompts included the same text of the task given to participants, previous submissions, personality prompts, current copy, elapsed time, history of its own actions, history of its chain-of-thought, chat history, general instructions, and a screenshot of the image after each change by the participant. We used OpenAI’s multimodal gpt-4o model (specifically gpt-4o-2024-08-06). Additional details in the AI agent are in SI Section  D.
Using Pairit, we designed the AI agent to be able to take the same actions that a human participant could take, apart from submitting ads. The actions include sending chat messages, editing any element of the ad copy (including the headline, primary text, and description), selecting images, generating images using the Dall-E 3 interface, or waiting (i.e., not taking any action). The agent is prompted every 10 seconds on whether to engage to take an action.
To ensure the agent is taking actions appropriately, we use chain-of-thought prompting using OpenAI’s structured output feature (wei2023chainofthought). Specifically, we prompt the model with questions to reflect on the state of the collaboration. These questions were necessary to avoid undesirable model behavior, such as repeatedly sending the same message.
For the randomization of AI personalities, we prompt the model with prompts that induce high or low levels of the Big Five personality traits. Each of the five traits is independently randomized to be high ($p=0.5$) or low ($p=0.5$; Figure 1B) using P2 prompting, which generates a detailed description of individuals with the traits (jiang2023evaluating). See SI Section E for covariate balance of the randomization of each personality prompt.
Procedure
As soon as participants were redirected to our platform from Prolific, they were randomized into either the human-human or the human-AI condition (Figure 1A). In the human-human condition, which we do not discuss in this study, participants joined a queue until another participant was available, at which point they were paired with each other. In the human-AI condition, which is the focus of this study, a participant joined a simulated queue, in which they waited for a random amount of time between 1 and 5 seconds, after which they were paired with an AI agent. We do not reveal whether or not the partner is a human or an AI until the post-task survey. Before the participants enter the task, they answered a 10-item survey, on a 7-point Likert scale, to measure their Big Five personality traits (rammstedt2007personality). See SI Section E for summary statistics and covariate balance of the human-AI randomization and SI Section G for the survey items.
Following the pairing and completion of the pre-task survey, participants entered a collaborative workspace where they could chat and collaboratively develop and submit display ads (Figure 1B). Participants were allotted 40 minutes to create and submit as many display ads as possible. After 40 minutes, participants were automatically re-directed to a post-task survey.
Upon completion of the task, participants completed a 35-item survey assessing teamwork quality and perception of AI (hoegl2001). The original instrument included 38 questions on a 7-point Likert scale, but three communication-related items were excluded as they were irrelevant to a one-time online collaborative effort. Next, the survey included two items on prior AI experience (i.e., “I have previously used AI chatbots (e.g., ChatGPT, Bard),” and “My experiences with AI chatbots have been positive”), one item on whether they perceived their partner as AI (i.e., “I thought my collaborator was an AI during the task”), and a final item disclosing the partner’s identity (human or AI) and asking if this altered their view of the collaboration quality (i.e., “Your collaborator was an AI/a human. How has this knowledge affected your perception of the collaboration’s quality?”). 85% of participants completed the post-task survey.
Participants
We pre-registered a target sample of 2,500 individuals recruited via Prolific (prolific.com) located in the US, with balanced representation by gender and ethnicity. Ultimately, 2310 participants completed the task. Of the 2500 individuals who entered, 23 failed to join the queue and were excluded from the study. Additionally, 167 participants either withdrew or exceeded the time limit prior to being paired with a collaborator. The overall attrition rate was 7.6%. The study was conducted from October 15 to 18, 2024, and took a median of 46.32 minutes to complete. The participants were paid \$9.
Incentives
To incentivize the production of high-quality ads, we informed the participants of their eligibility for additional \$100 as bonus payment, determined by the volume and quality of their submissions. Participants were explicitly instructed that “the greater the number of ads, the greater your chances—but not if the ads are of low quality." Ultimately, two participants were awarded \$100 each.
Human evaluation of ad quality
To obtain human ratings of ad quality, we recruited a separate sample of 1,995 participants from Prolific based in the US, with representative stratification across gender, age, and ethnicity. A total of 1,200 individuals entered the survey. Of these participants, 5 dropped out before submitting their surveys, resulting in a dropout rate of 0.42%. We built a custom platform for this survey, run on Google Cloud Platform’s App Engine. The code is available on GitHub. This evaluation was conducted from November 7 to 9, 2024, and took a median of 16.47 minutes to complete.
To obtain ratings for all ads while avoiding survey fatigue, we created a random sample, with random order, of 40 ads per participant. To ensure that each ad received at least 3 ratings, we produced a set of 1,300 samples. As a participant entered our survey platform, we drew one sample from the set, without replacement, to provide to the participant, with each participant receiving a unique sample.
To estimate ad performance, we created and used mockups of display ads in the survey (see SI Section C). For each ad, we asked participants to provide three ratings regarding the quality of the text, the image, and the estimated click through rate. Each item was rated on a 7-point Likert scale. The first rating was based on the following prompt “The text is present, clear, relevant, and engaging"; the second on “The image is visually appealing"; and the third on “I am likely to click on this ad."
Field evaluation of ad performance
We evaluated real-world ad performance on the X platform, pre-registered at osf.io/95dhu. Key metrics included click-through rate (CTR), cost-per-click (CPC), view-through rate (VTR; fraction of document viewed), and view-through duration (VTD; seconds). VTR and VTD were tracked via a unique DocSend link for each ad. Our ad campaigns assessed causal ad effects conditional on online algorithms and divergent targeting (braun2024adplatforms, braun2024abtesting). No holdout group was used, as unique DocSend links imply zero visits without ads (braun2024adplatforms).
We sampled 2,000 ads from 11,138, selecting 1–2 ads per 1,834 human-human and human-AI teams using stratified sampling. Ads were grouped by team type and divided into 10 strata based on human-predicted click likelihood. Eight ads were excluded for potential policy violations (e.g., violence, sexual, or drug content). To avoid audience overlap, 400 campaigns each ran five unique ads as split tests. Each campaign targeted 133 random ZIP codes with population between 10,000 and 100,000, as per the 2020 American Community Survey. One-way ANOVA yielded non-significant results across ZIP codes: $F(399,53199)=0.954$, $p=0.734$ for population; $F(399,53199)=0.973$, $p=0.636$ for income. Ads were run from January 21 to February 9, 2025, with 50 campaigns active for two days each due to platform constraints. Campaign random effects controlled for temporal confounders. In our regression analyses, advertising spend was controlled for to account for auto-bidding.
Model specifications
To evaluate the effect of personality trait interactions on human-AI team outcomes, we use the following regression model:
$Y_{J}=\sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}+\epsilon_{i}\text{,}$ (1)
where $Y_{J}$ is the outcome for observation $i$, BF are the Big Five traits, $h\in\text{BF}$ is a human trait, $a\in\text{BF}$ is a prompted AI trait, $\beta_{h}$ and $\beta_{a}$ are coefficients for the main effects of human and AI traits respectively, $\beta_{ha}$ captures the interaction between human trait $h$ and AI trait $a$, $T_{h,i}$ is the normalized (0 to 1) score on trait $h$ for the human participant associated with observation $i$, $T_{a,i}$ is an indicator equal to 1 if the AI partner was prompted to exhibit high levels of trait $a$ (0 for low), and $\epsilon_{i}$ is the error term. The unit of observation $i$ varies by outcome: for ad quality ratings, it is the individual rating (with standard errors clustered at the ad level); for productivity, teamwork quality measures, and message type fractions, it is the individual participant (with heteroskedasticity-robust standard errors). The outcomes of interest $Y_{i}$ include human-evaluated ad quality, productivity (i.e., the number of ads submitted), and teamwork quality measures.
To evaluate the field experiment, we use the following mixed effects model:
$Y_{ic}=$ $\ \sum_{h\in\text{BF}}\beta_{h}T_{h,i}+\sum_{a\in\text{BF}}\beta_{a}T_{a,i}+\sum_{h\in\text{BF}}\sum_{a\in\text{BF}}\beta_{ha}T_{h,i}T_{a,i}$
$+\text{Text}_{ic}+\text{Image}_{ic}+\text{Click}_{ic}+\text{Spend}_{ic}+u_{c}+\epsilon_{ic}\text{,}$ (2)
where $Y_{ic}$ represents the outcome (e.g., CPC, CTR, VTR, VTD) for ad $i$ in campaign $c$, Image, Text, and Click are human-rated quality measures, $\text{Spend}_{ic}$ is the campaign spend, $u_{c}\sim N(0,\sigma^{2}_{u})$ is the random intercept for campaign $c$, and the rest are the same as in Equation 1.
Acknowledgments
We thank Dean Eckles, John Horton and members of the MIT Social Analytics Lab for their invaluable discussions. This work was supported by the MIT Initiative on the Digital Economy (IDE) at the MIT Sloan School of Management and was approved by the Massachusetts Institute of Technology Institutional Review Board (IRB).